# Investigating Domain Gaps for Indoor 3D Object Detection

### v1

@@ Abstract.
As a fundamental task for indoor scene understanding, 3D object detection has been extensively studied, and the accuracy on indoor point cloud data has been substantially improved. However, existing researches have been conducted on limited datasets, where the training and testing sets share the same distribution. In this paper, we consider the task of adapting indoor 3D object detectors from one dataset to another, presenting a comprehensive benchmark with ScanNet, SUN RGB-D and 3D Front datasets, as well as our newly proposed large-scale datasets ProcTHOR-OD and ProcFront generated by a 3D simulator. Since indoor point cloud datasets are collected and constructed in different ways, the object detectors are likely to overfit to specific factors within each dataset, such as point cloud quality, bounding box layout and instance features. We conduct experiments across datasets on different adaptation scenarios including synthetic-to-real adaptation, point cloud quality adaptation, layout adaptation and instance adaptation, analyzing the impact of different domain gaps on 3D object detectors. We also introduce several approaches to improve adaptation performances, providing baselines for domain adaptive indoor 3D object detection, hoping that future works may propose detectors with stronger generalization ability across domains. Our benchmark datasets and baseline code can be found in: https://github.com/JeremyZhao1998/DAVoteNet-release.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758275isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Object detectionccs: Computing methodologies Scene understanding
@@ 1. Introduction
Only a few studies (e.g., (Yunsong et al., 2024)) have explored domain adaptation techniques for indoor 3D object detection, while their focus has been limited to the synthetic-to-real adaptation scenario. However, indoor 3D object detection encompasses a broader range of domain variations and practical application scenarios, which are not adequately represented or evaluated by current benchmarks. To facilitate a more systematic investigation of the various sources of domain shift that affect the performance of indoor 3D object detectors, we introduce a suite of comprehensive benchmarks. These benchmarks are constructed from both widely used public datasets and newly developed datasets designed to capture diverse conditions.
@@
After selecting a consistent subset of instance categories across datasets, we combine the five datasets to construct a suite of domain adaptation benchmarks, covering the following scenarios: (1) Synthetic-to-real adaptation: Following (Yunsong et al., 2024), we evaluate 3D Front $\rightarrow$ ScanNet and 3D Front $\rightarrow$ SUN RGB-D, and additionally propose ProcTHOR-OD $\rightarrow$ ScanNet as a new synthetic-to-real setting where the source domain has larger data scale. (2) Point cloud quality adaptation: We test ScanNet $\rightarrow$ SUN RGB-D to assess the performance drop when transferring from high-quality scans to lower-quality, single-frame point clouds. (3) Layout adaptation: To isolate the impact of layout differences, we evaluate ProcFront $\rightarrow$ 3D Front. The two datasets share the same instance models but differ in scene layouts. (4) Instance adaptation: To examine domain gaps at the instance level, we evaluate ProcTHOR-OD $\rightarrow$ ProcFront, where the layouts remain the same but object appearances vary.
Under such benchmarks, we conduct experiments on the commonly used VoteNet (Qi et al., 2019) detector. As shown in fig. 1(b), the detector performs well within each dataset (blue bars, trained and evaluated within target dataset), but cross-dataset evaluation shows a significant performance drop (orange bars, trained by source and evaluated on target dataset). We also implement several commonly used domain adaptation approaches in other tasks to improve adaptation performances. These approaches serve as baselines for domain adaptive indoor 3D object detection, hoping that future works may propose detectors or frameworks with stronger generalization ability across domains. Our benchmark datasets and baseline code can be found in: https://github.com/JeremyZhao1998/DAVoteNet-release.
@@ 2.1. Indoor point cloud datasets
Despite their contributions to downstream tasks, existing datasets and benchmarks remain limited in scale, quality, and label diversity. They also assume identical distributions between training and testing sets, whereas this paper focuses on cross-domain adaptation for indoor 3D object detection. To this end, we introduce the ProcTHOR-OD and ProcFront datasets using the AI2-THOR platform (Deitke et al., 2022), and combine them with existing datasets to establish new domain adaptation benchmarks.
@@ 2.3. Domain adaptation for point clouds
Prior work on 3D domain adaptation has primarily focused on classification, segmentation, and LiDAR-based detection tasks. PointDA (Qin et al., 2019) first introduced a domain adaptation method for 3D object classification by reconstructing local point features. Subsequent studies (Shen et al., 2022; Cardace et al., 2023) proposed global feature reconstruction and self-training strategies to improve adaptation performance. For semantic segmentation, (Ding et al., 2022) explored synthetic-to-real adaptation via data augmentation and mixing.
LiDAR-based detection in autonomous driving scenes has a close relationship with indoor point cloud object detection. The pioneer work (Wang et al., 2020) evaluate cross-dataset performance of LiDAR detectors, followed by multiple techniques (Hu et al., 2023; Liu et al., 2024; Chang et al., 2024) that enhance domain adaptation performance by self-training, point completion, and pseudo-label refinement. Nonetheless, adapting to indoor 3D detection poses distinct challenges: (1) unlike audonomous driving datasets where all point clouds are LiDAR-generated, indoor point clouds arise from diverse sources, leading to more complex and varied domain gaps; (2) LiDAR point cloud detection typically focuses on a single object category (e.g., cars), whereas indoor detection involves a wider range of object classes (more than 10) with richer semantics. In this paper, we investigate the domain gaps for indoor 3D object detection task, providing the first comprehensive benchmark and several baselines.
@@ 3.3.1. Analysis of dataset status
In fig. 2 we present typical scenes of each dataset. In table 1 and fig. 3 we show the statistical information of the datasets. We analyze the differences of these datasets as follows: (1) Data scale: As shown in table 1, while existing real-world datasets are limited in scale and annotation quality, our proposed synthetic datasets offer low-cost, high-precision annotations with 10k scenes and 238k bounding boxes—an order of magnitude more than their real-world counterparts. (2) Style(synthetic to real): As illustrated in fig. 2, the style difference exists between synthetic and real-world datasets. Though providing high quality point clouds, synthetic datasets lack the texture details and material realism of objects in real-world environments. (3) Point cloud quality: As can be visualized in fig. 2, thoroughly scanned ScanNet dataset have high quality while SUN RGB-D point clouds exhibit obvious point omission. The synthetic datasets have high quality point clouds without the problem of scan omission. (4) Room layout: As shown in table 1, ScanNet and SUN RGB-D are collected from real-world indoor environments, featuring naturally cluttered object arrangements, whereas 3D Front consists of expert-designed scenes with clean and organized layouts. In contrast, our proposed ProcTHOR dataset adopts automatically generated layouts, which differ from human-designed ones in style but offer significantly greater scale and diversity. (5) Instance features: Different datasets exhibit distinct instance-level characteristics, even within the same object category. As shown in fig. 2, objects such as beds can vary significantly in appearance. Moreover, as illustrated in fig. 3, the distribution of instance counts also differs notably, with our synthetic dataset providing a substantially larger number of instances.
@@ 3.3.2. Domain adaptation scenarios
Synthetic to real adaptation: Fully leveraging the large-scale, high-quality synthetic data, the detectors should be capable of adapting to real-world scenarios to meet practical requirements. Under this setting, we use our proposed ProcTHOR-OD as source domain and ScanNet as target domain. Following (Yunsong et al., 2024), we also evaluate 3D Front $\rightarrow$ ScanNet and 3D Front $\rightarrow$ SUN RGB-D.
@@ 4.1. Setup and details
[…]
During inference, we post-process the proposals with dropping empty boxes (less than 5 points in a box), dropping too little boxes (boxes with side length less than 0.001 m) and 3D NMS (non-maximum suppression) with IoU threshold of 0.25. We report the evaluation metric of mAP (mean average precision) following the protocal of (Song and Xiao, 2016) with IoU threshold 0.25. More details can be found in our link of benchmark datasets and code.
@@ 4.2. Analysis of domain gaps
UDA VSS (Ding et al., 2022) 20.31 16.72 10.61 25.26 28.06 24.32
MT (Tarvainen and Valpola, 2017) 20.43 15.03 10.32 25.25 31.77 27.43
RV (Fan et al., 2022) 20.89 15.27 10.40 24.92 32.45 28.50
@@ 4.3.1. Using target domain priors
Using object size prior: Different datasets exhibit average object size difference. By simply adopting the target domain statistical prior of mean sizes for each category instead of precise labels, reported in line “Size”, the detector shows performance gains in most benchmarks where the source and target domains have a larger instance gap. (In “pf2front” and “proc2pf” where the source and target domains share a large number of instances, the size prior seems to have no effect.) This observation shed a light on future studies to estimate target domain object sizes.
@@ 4.3.2. Unsupervised Domain Adaptation
As shown in the “UDA” block of table 8, VSS augmentation performs well in synthetic-to-real settings but lacks generality for layout and instance adaptation. (The augmented data has lower point cloud quality, thus harms the performance where the target domain has high-quality point clouds in “pf2front” and “proc2pf”.) The mean-teacher approach consistently works across all baselines, demonstrating its generality, and reliable voting (RV) consistently improves mean-teacher results, indicating the effectiveness of pseudo-label filtering. More details can be seen in our baseline code link.

### v2

@@ Abstract.
As a fundamental task for indoor scene understanding, 3D object detection has been extensively studied, and the accuracy on indoor point cloud data has been substantially improved. However, existing researches have been conducted on limited datasets, where the training and testing sets share the same distribution. In this paper, we consider the task of adapting indoor 3D object detectors from one dataset to another, presenting a comprehensive benchmark with ScanNet, SUN RGB-D and 3D Front datasets, as well as our newly proposed large-scale datasets ProcTHOR-OD and ProcFront generated by a 3D simulator. Since indoor point cloud datasets are collected and constructed in different ways, the object detectors are likely to overfit to specific factors within each dataset, such as point cloud quality, bounding box layout and instance features. We conduct experiments across datasets on different adaptation scenarios including synthetic-to-real adaptation, point cloud quality adaptation, layout adaptation and instance adaptation, analyzing the impact of different domain gaps on 3D object detectors. We also introduce several approaches to improve adaptation performances, providing baselines for domain adaptive indoor 3D object detection, hoping that future works may propose detectors with stronger generalization ability across domains. The benchmark datasets and baseline code are available on our project homepage: https://jeremyzhao1998.github.io/DAVoteNet-release.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758275isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Object detectionccs: Computing methodologies Scene understanding
@@ 1. Introduction
Only a few studies (e.g., (Wang et al., 2024)) have explored domain adaptation techniques for indoor 3D object detection, while their focus has been limited to the synthetic-to-real adaptation scenario. However, indoor 3D object detection encompasses a broader range of domain variations and practical application scenarios, which are not adequately represented or evaluated by current benchmarks. To facilitate a more systematic investigation of the various sources of domain shift that affect the performance of indoor 3D object detectors, we introduce a suite of comprehensive benchmarks. These benchmarks are constructed from both widely used public datasets and newly developed datasets designed to capture diverse conditions.
@@
After selecting a consistent subset of instance categories across datasets, we combine the five datasets to construct a suite of domain adaptation benchmarks, covering the following scenarios: (1) Synthetic-to-real adaptation: Following (Wang et al., 2024), we evaluate 3D Front $\rightarrow$ ScanNet and 3D Front $\rightarrow$ SUN RGB-D, and additionally propose ProcTHOR-OD $\rightarrow$ ScanNet as a new synthetic-to-real setting where the source domain has larger data scale. (2) Point cloud quality adaptation: We test ScanNet $\rightarrow$ SUN RGB-D to assess the performance drop when transferring from high-quality scans to lower-quality, single-frame point clouds. (3) Layout adaptation: To isolate the impact of layout differences, we evaluate ProcFront $\rightarrow$ 3D Front. The two datasets share the same instance models but differ in scene layouts. (4) Instance adaptation: To examine domain gaps at the instance level, we evaluate ProcTHOR-OD $\rightarrow$ ProcFront, where the layouts remain the same but object appearances vary.
Under such benchmarks, we conduct experiments on the commonly used VoteNet (Qi et al., 2019) detector. As shown in fig. 1(b), the detector performs well within each dataset (blue bars, trained and evaluated within target dataset), but cross-dataset evaluation shows a significant performance drop (orange bars, trained by source and evaluated on target dataset). We also implement several commonly used domain adaptation approaches in other tasks to improve adaptation performances. These approaches serve as baselines for domain adaptive indoor 3D object detection, hoping that future works may propose detectors or frameworks with stronger generalization ability across domains. Our project homepage is: https://jeremyzhao1998.github.io/DAVoteNet-release.
@@ 2.1. Indoor point cloud datasets
Despite their contributions to 3D foundation model training(Yang et al., 2024; Wang et al., 2025b) and downstream tasks including 3D grounding(Guo et al., 2025; Li et al., 2025b, a), VQA(Mo and Liu, 2024; Luo et al., 2025b) and 3D content generation(Qin et al., 2025; Ye et al., 2024a, b; Deng et al., 2025a; Liang et al., 2025), existing datasets remain limited in scale, quality, and label diversity. They also assume identical distributions between training and testing sets, whereas this paper focuses on domain adaptation for indoor 3D object detection. To this end, we introduce synthetic ProcTHOR-OD and ProcFront datasets, and combine them with existing datasets to establish new domain adaptation benchmarks.
@@ 2.3. Domain adaptation for point clouds
Domain adaptation has been extensively studied in 2D computer vision (Pei et al., 2024; Deng et al., 2025b; Chen et al., 2018; Pei et al., 2023; Zhang et al., 2025; Xu et al., 2022; Chen and Liu, 2020) and downstream tasks such as detection(Zhao et al., 2023; Pan et al., 2025), segmentation(Luo et al., 2025d; Wang et al., 2025a) and retrieval(Chen et al., 2021; Liu et al., 2021; Luo et al., 2025a, c). Prior works on 3D domain adaptation has primarily focused on classification, segmentation, and LiDAR point cloud understanding. PointDA (Qin et al., 2019) first introduced a domain adaptation method for 3D classification by reconstructing local point features. Subsequent studies (Shen et al., 2022; Cardace et al., 2023) proposed global feature reconstruction and self-training to improve adaptation performance. For semantic segmentation, (Ding et al., 2022) explored synthetic-to-real adaptation via data augmentation and mixing.
LiDAR-based detection in autonomous driving scenes has a close relationship with indoor point cloud object detection. The pioneer work (Wang et al., 2020) evaluate cross-dataset performance of LiDAR detectors, followed by multiple techniques (Hu et al., 2023; Liu et al., 2024; Chang et al., 2024) that enhance domain adaptation performance by self-training, point completion, and pseudo-label refinement. Nonetheless, indoor 3D detection poses distinct challenges: (1) unlike audonomous driving datasets that are all LiDAR-generated, indoor point clouds arise from diverse sources, leading to more complex and varied domain gaps; (2) LiDAR point cloud detection typically focuses on the car category, whereas indoor detection involves a wider range of object classes (more than 10) with richer semantics. In this paper, we investigate the domain gaps for indoor 3D object detection task, providing the first comprehensive benchmark and several baselines.
@@ 3.3.1. Analysis of dataset status
In fig. 2 we present typical scenes of each dataset. In table 1 and fig. 3 we show the statistical information of the datasets. We analyze the differences of these datasets as follows: (1) Data scale: As shown in table 1, while existing real-world datasets are limited in scale and annotation quality, our proposed synthetic datasets offer low-cost, high-precision annotations with 10k scenes and 238k bounding boxes which is an order of magnitude more than ScanNet. (2) Style(synthetic to real): As illustrated in fig. 2, the style difference exists between synthetic and real-world datasets. Though providing high quality point clouds, synthetic datasets lack the texture details and material realism in real-world environments. (3) Point cloud quality: As shown in fig. 2, ScanNet dataset have higher quality point clouds than SUN RGB-D which exhibits obvious point omission. Point clouds in synthetic datasets are randomly sampled without scan omission. (4) Room layout: As shown in table 1, ScanNet and SUN RGB-D are collected from real-world indoor environments, featuring naturally cluttered object arrangements, whereas 3D Front consists of expert-designed scenes with clean and organized layouts. In contrast, our proposed ProcTHOR dataset adopts automatically generated layouts, which differ from human-designed ones in style but offer significantly greater scale and diversity. (5) Instance features: Different datasets exhibit distinct instance-level characteristics, even within the same object category. As shown in fig. 2, objects such as beds can vary significantly in appearance, which may mislead the detector to produce wrong prediction on categories. Moreover, as illustrated in fig. 3, the distribution of instance counts also differs notably, with our synthetic dataset providing a substantially larger number of instances. The instance number distribution may also influence the prediction of cross-domain 3D detectors.
@@ 3.3.2. Domain adaptation scenarios
Synthetic to real adaptation: Fully leveraging the large-scale, high-quality synthetic data, the detectors should be capable of adapting to real-world scenarios to meet practical requirements. Under this setting, we use our proposed ProcTHOR-OD as source domain and ScanNet as target domain. Following (Wang et al., 2024), we also evaluate 3D Front $\rightarrow$ ScanNet and 3D Front $\rightarrow$ SUN RGB-D.
@@ 4.1. Setup and details
In our experiments, we apply commonly used indoor 3D object detector VoteNet (Qi et al., 2019) to evaluate the cross domain performance and domain adaptation approaches. We randomly sample 40,000 points per scene for training. In addition to coordinates, we include height and RGB color features for each point. We train the network end-to-end with AdamW optimizer and batch size 32 for 90 epochs.
During inference, we post-process the proposals with dropping empty boxes (less than 5 points in a box), dropping too little boxes (boxes with side length less than 0.001 m) and 3D NMS (non-maximum suppression) with IoU threshold of 0.25. We report the evaluation metric of mAP (mean average precision) following the protocal of (Song and Xiao, 2016) with IoU threshold 0.25. More details can be found in our open source code.
@@ 4.2. Analysis of domain gaps
UDA VSS (Ding et al., 2022) 20.31 16.72 10.61 25.26 28.15 24.32
MT (Tarvainen and Valpola, 2017) 20.43 15.03 10.32 25.25 32.70 27.44
RV (Fan et al., 2022) 20.89 15.27 10.40 24.92 32.74 27.78
@@ 4.3.1. Using target domain priors
Using object size prior: Different datasets exhibit average object size difference. By simply adopting the target domain statistical prior of mean sizes (instead of precise labels) for each category, reported in line “Size”, the detector shows performance gains across all the benchmarks. This observation shed a light on future studies to estimate target domain object sizes.
@@ 4.3.2. Unsupervised Domain Adaptation
As shown in the “UDA” block of table 8, VSS augmentation performs well in synthetic-to-real settings but lacks generality for layout and instance adaptation. (The augmented data has lower point cloud quality, thus harms the performance where the target domain has high-quality point clouds in “pf2front” and “proc2pf”.) The mean-teacher approach consistently works across all baselines, demonstrating its generality, and reliable voting (RV) consistently improves mean-teacher results, indicating the effectiveness of pseudo-label filtering. More details can be seen in our open-source code.